OpenStudy (anonymous):

the intergral of (4/(3-6x))dx

OpenStudy (anonymous):

\[-\frac23\ln|3-6x|+c\]

OpenStudy (saifoo.khan):

Oops, that's a 2.

OpenStudy (saifoo.khan):

\[\int\limits \frac{4}{3(1-2x)}\]

OpenStudy (anonymous):

\[\int\limits \frac{f'(x)}{f(x)}dx=\ln f(x)+c\]

OpenStudy (callisto):

I felt very stupid to have done that :S

OpenStudy (callisto):